There are hundreds of apples on the trees, so you randomly choose just 83 apples and get a mean weight of 87.53 grams with a standard deviation of 6.7 grams. Assuming that the weight of the apples is normally distributed, what is the standard error of the 90% confidence interval for the mean weight of the apples.


1
Expert's answer
2022-03-29T15:53:27-0400

"SE=\frac{s}{\sqrt{n}}=\frac{6.7}{\sqrt{83}}=0.74."